OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–1, 4) and is parallel to the line y = 2x – 3? OPTIONS: A) x – 2y = 6 B) x + 2y = 6 C) 2x – y = –6 D) 2x + y = –6

OpenStudy (helder_edwin):

do u know what is the slope of the line y=2x-3 ?

OpenStudy (anonymous):

2

OpenStudy (helder_edwin):

right

OpenStudy (anonymous):

So, (1,4) is the point we given. TO eb parrallel, the slopes have to be the same. Using the form y=mx+b, where m is the slope, we can say the new equation has to be y=2x+b so using the (1,4), we go 4=2(1)+b, solve for b b=2 so the equation of the line is y=2x+2

OpenStudy (anonymous):

but thats not one of my options

OpenStudy (anonymous):

Woops, I missed a negative sign

OpenStudy (anonymous):

b=6, so y=2x+6

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

but wich option is it?

OpenStudy (anonymous):

Do you understand the process though?

OpenStudy (anonymous):

Just rearrange the equation.

OpenStudy (anonymous):

is it D? @dannym

OpenStudy (anonymous):

or B @dannym